OpenStudy (anonymous):

What is the solution to the equation? 7(5 + 9) = 7 • x + 7 • 9 A. x = 1 B. x = 5 C. x = 7 D. x = 9

OpenStudy (anonymous):

i bet you can guess it

OpenStudy (anonymous):

it is the only number missing

OpenStudy (anonymous):

5

OpenStudy (anonymous):

right? @satellite73

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

I got 2 more questions

OpenStudy (anonymous):

\[7(5+9)=7\times 5+7\times 9\] is the distributive property in action

OpenStudy (anonymous):

aidian can i have best response?

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

I'm gonna close this question now

OpenStudy (anonymous):

bye bye